OpenStudy (saylilbaby):

a cylindrical jar has a diameter of 9 in and height of 12 in. The water surface is 2.5 in from the top. About how much water is in the jar? use 3.14 for n.

OpenStudy (saylilbaby):

@Ria23

OpenStudy (dan815):

thats not for n that is for pi

OpenStudy (dan815):

bu nice lol.. n does look close to \[\pi\]

OpenStudy (saylilbaby):

what ???? that is the question @dan815 i need help solving it

OpenStudy (dan815):

what do u have so far

OpenStudy (saylilbaby):

nun @dan815

OpenStudy (dan815):

okay start with a drawing

OpenStudy (saylilbaby):

@TheSmartOne

OpenStudy (saylilbaby):

@haleyelizabeth2017

OpenStudy (dan815):

do some work

OpenStudy (dan815):



Directrix (directrix):

Because the base is a circle, the formula can be written as: V = pi * r^2 * h The height of the cylinder is 12 but the water volume which you are asked to find has a height of (12 - 2.5). So, you will need to crank out: V = pi * (4.5)^2 * 9.5 and approximate pi by 3.14. Note that the base radius is 4.5 because the radius of a circle is one-half of its diameter. @Saylilbaby